Amelia Enterprises has presented the following information for the past three months operations:

Month Units Average costs

June 2,400 $10.00

July 4,800 $6.00

August 6,000 $5.20
a. Using the high-low method, calculate the fixed cost per month and variable cost per unit.
b. What would total costs be for a month with 5,000 units produced?
Answer #1

Answer

a) Highest number of units produced

= 6000 units@$5.20

= $31,200


Lowest number of units produced

= 2400 units @$10

= $24000


Variable cost per unit using high low method will be
Change in cost/Change in units produced
= 31200-24000/6000-2400
= 7200/3600

= $2 per unit
Variable cost per unit

= $2 per unit
Fixed cost= $31200-(6000*2) = $19200

b) Total cost for 5000 units produced:

Variable cost= 5000*$2= $10000
Fixed cost = $19200
Total cost= $29200

Please like